# A Low-Rank ADMM Splitting Approach for Semidefinite Programming

### 14 Mar 2024 v1

@@ Abstract
We introduce a new first-order method for solving general semidefinite programming problems, based on the alternating direction method of multipliers (ADMM) and a matrix-splitting technique. Our algorithm has an advantage over the Burer-Monteiro approach as it only involves much easier quadratically regularized subproblems in each iteration. For a linear objective, the subproblems are well-conditioned quadratic programs that can be efficiently solved by the standard conjugate gradient method. We show that the ADMM algorithm achieves sublinear or linear convergence rates to the KKT solutions under different conditions. Building on this theoretical development, we present LoRADS, a new solver for linear SDP based on the Low-Rank ADMM Splitting approach. LoRADS incorporates several strategies that significantly increase its efficiency. Firstly, it initiates with a warm-start phase that uses the Burer-Monteiro approach. Moreover, motivated by the SDP low-rank theory (So et al. 2008), LoRADS chooses an initial rank of logarithmic order and then employs a dynamic approach to increase the rank. Numerical experiments indicate that LoRADS exhibits promising performance on large-scale SDP problems. A noteworthy achievement of LoRADS is its successful solving of a matrix completion problem with $15,694,167$ constraints and a matrix variable of size $000\times 40,$ 000 × 40 , 000 in $351$ seconds.
@@ 1 Introduction
SDPs have a wide range of applications in practice, such as the MaxCut problem (Goemans and Williamson, 1995), optimal power flow (Lavaei and Low, 2011), combinatorial optimization (Boyd and Vandenberghe, 1997) and sensor network localization (So and Ye, 2007), to name a few.
Traditionally, interior point methods are the go-to solution for solving SDP problems. It has been well known that for SDP, interior point methods can achieve high accuracy in polynomial time. See  Alizadeh (1995); Nesterov and Nemirovskii (1994); Alizadeh (1991). Despite their effectiveness, traditional interior point methods struggle with scalability for large-scale problems due to the necessity of solving Newton systems repeatedly, which is computationally intensive. To bypass the Newton system and enhance scalability, more recent research has shifted towards the first-order methods (Wen et al., 2010; Lan et al., 2011; O’donoghue et al., 2016). A significant area of interest has been the use of penalty methods, such as the augmented Lagrangian method and alternating direction method of multipliers (ADMM), for solving SDPs (Wen et al., 2010; Zhao et al., 2010; Yang et al., 2015; Wang et al., 2023). Specifically, Zhao et al. (2010) introduced SDPNAL, which employs the augmented Lagrangian method on dual SDP and uses an inexact semi-smooth Newton-CG method to solve the inner subproblem. The subsequent work (Yang et al., 2015) presented an enhanced algorithm SDPNAL+, which achieves improved efficiency and robustness and adequately addresses the numerical difficulty of SDPNAL when solving degenerate SDPs. O’donoghue et al. (2016) proposes a general splitting conic solver (SCS) for linear and conic programming, which applies ADMM to solve the homogeneous self-dual embedding associated with SDP. For large-scale SDP, a critical issue is the quadratically growing ($\mathcal{O}(n^{2})$) memory cost in maintaining the matrix variable. For a class of rank-constrained matrix problems, Yurtsever et al. (2017) apply the conditional gradient method to their convex relaxation problems. To reduce the storage overheads, they employ random sketching to keep track of a low-rank approximation of the matrix variable. In a follow-up work by Yurtsever et al. (2021), random sketching is further incorporated in the augmented Lagrangian-based method to deal with more general trace-constrained SDPs.
@@
Burer and Monteiro (2003, 2005) propose an effective limited memory BFGS augmented Lagrangian algorithm to handle the low-rank formulation. Another popular low-rank approach for SDP is the Riemannian optimization methods (Journée et al., 2010; Wen and Yin, 2013; Boumal, 2015). These methods are built upon the assumption that the feasible set of (3) is a Riemannian manifold. Within the low-rank formulation, Tang and Toh (2023) proposes a rank-support adaptive Riemannian optimization method for solving the general convex SDP problems and gives the convergence rate analysis. There are also several different ways to solve the low-rank formulation of SDP. Bellavia et al. (2021) propose a relaxed variant of the interior point method, which exploits the low-rank structure of the optimal solution. Erdogdu et al. (2022) apply the block-coordinate ascent method to solve the low-rank reformulation of SDP with diagonal constraints and establish global convergence towards the first-order stationary point. In the BM formulation, a critical step is to determine the rank of the matrix variable. The rank $r$ is chosen to ensure that either an optimal solution for (3) is attainable or a satisfactory approximate low-rank solution is obtained. In particular, So et al. (2008) demonstrate the existence of a solution of logarithmic rank order that still maintains good approximation property. This discovery forms the basis of an empirical tactic to initiate the algorithm with a solution of rank $\mathcal{O}(\log(m))$, which will be employed in our approach.
@@
Our paper’s contributions are summarized as follows: First, we introduce a novel ADMM algorithm tailored for general SDPs that accommodate nonlinear and convex objectives. By introducing a new matrix splitting approach, we recast the SDP into a bilinearly constrained and quadratically penalized nonconvex problem, much different from the Burer-Monteiro approach. This formulation offers two primary benefits for ADMM implementation: 1) For a linear objective, both the alternating steps in ADMM result in a quadratic subproblem associated with a positive definite matrix, hence, can be efficiently solved by either direct methods or indirect ones such as the conjugate gradient descent method for a well-conditioned system enabled by the splitting technique. This simplification provides a substantial advantage over the nonconvex quartic subproblems inherent in the BM formulation; 2) The split searching approach of ADMM addresses some challenges that traditional factorization methods cannot solve. Under mild conditions, we prove that the ADMM algorithm achieves a sublinear rate of convergence to the KKT solutions. Additionally, when the Łojasiewicz inequality is satisfied with an exponent of ${1}/{2}$, we enhance the convergence result to a linear rate, offering a significant improvement in the algorithm’s efficiency.
Second, we present a new SDP solver, which employs low-rank factorization and ADMM through matrix splitting (LoRADS). This solver integrates the proposed ADMM along with several new implementation strategies that significantly improve the practical efficiency. Our empirical evaluations highlight two critical heuristics: 1) aggressive rank reduction to logarithmic order with a dynamic rank selection strategy ensuring the stability, and 2) a warm-start procedure employing the Burer-Monteiro algorithm.
@@
By leveraging these enhancements, we compare LoRADS with the leading SDP solvers on a wide range of problems, including MaxCut, matrix completion, and general SDP benchmarks in SDPLIB and Mittlemann’s dataset. Extensive numerical studies demonstrate that LoRADS is highly competitive and often outperforms the popular SDP solvers. Furthermore, LoRADS demonstrates notable efficacy in addressing super large-scale problems, an area where current SDP solvers show limitations. A case in point is LoRADS’s ability to solve a matrix completion problem with $15,694,167$ constraints and a $40000\times 40000$ matrix variable within approximately 350 seconds.
@@
[…]
The remainder of our paper is organized as follows. Section 2 introduces the proposed ADMM algorithm with matrix splitting for semidefinite programming. Section 3 establishes the convergence and convergence rate of the proposed algorithm. Section 4 develops an enhanced solving approach based on ADMM and several useful implementation heuristics. Section 5 examines the empirical performance of our solver on various SDP problems. Finally, Section 6 concludes our paper and points out our future direction.
@@ 2.1 Bilinear Reformulations and The Penalized Problem
To address the challenges posed by large-scale SDP instances and to thoroughly explore the low-rank domain, we propose a bilinear factorization technique for problem (1), yielding the following formulation:
@@
To elucidate the relationship between model (4) and (6) with a suitably large parameter $\gamma$, it is imperative to introduce the following assumption.
@@
where $\nabla^{2}f(\cdot)[H]$ denotes the tensor product. From the symmetry of $f$, it is easy to verify that the partial gradients satisfy
$\frac{\partial{f(UV^{\top})}}{\partial U}=\nabla f(UV^{\top})V\quad{\rm and}\quad\frac{{f(UV^{\top})}}{\partial V}=\nabla f(VU^{\top})U.$
@@
$\begin{split}\Big{\|}2\nabla f(U^{*}{(U^{*})}^{\top})U^{*}+\sum_{i=1}^{m}\lambda_{i}^{*}A_{i}U^{*}\Big{\|}\leq\epsilon,\\ \big{\|}\mathcal{A}(U^{*}{(U^{*})}^{\top})-{b}\big{\|}\leq\epsilon,\end{split}$ (7)
@@
$\begin{split}\Big{\|}\nabla f(\overline{U}\overline{V}^{\top})\overline{V}+\gamma(\overline{U}-\overline{V})+\sum_{i=1}^{m}\overline{\lambda}_{i}A_{i}\overline{V}\Big{\|}\leq\epsilon,\\ \Big{\|}\nabla f(\overline{V}\overline{U}^{\top})\overline{U}+\gamma(\overline{V}-\overline{U})+\sum_{i=1}^{m}\overline{\lambda}_{i}A_{i}\overline{U}\Big{\|}\leq\epsilon,\\ \big{\|}\mathcal{A}(\overline{U}\overline{V}^{\top})-{b}\big{\|}\leq\epsilon.\end{split}$ (8)
@@
Let $(\overline{U},\overline{V},\overline{{\lambda}})$ be any $\epsilon$-KKT point of problem (6) in $\mathcal{B}_{\sigma}\big{(}0^{n\times r},0^{n\times r},0^{m}\big{)}$ and $\widehat{U}=\frac{\overline{U}+\overline{V}}{2}$. Suppose assumption 1 holds, then the following statements are valid.
  • (i)
    $\begin{split}&\left\|{\cal A}(\widehat{U}{\widehat{U}}^{\top})-b\right\|_{2}\leq\epsilon+\frac{\|\mathcal{A}\|_{2}(\epsilon+L_{f}^{\sigma}\sigma+\sqrt{s_{A}}\sigma^{2})^{2}}{4\gamma^{2}},\\ &\left\|2\nabla{f(\widehat{U}\widehat{U}^{\top})}\widehat{U}+2\sum_{i=1}^{m}\overline{\lambda}_{i}A_{i}\widehat{U}\right\|_{F}\leq 2\epsilon+{(2L_{0}^{\sigma}\sigma^{3}+3L^{\sigma}\sigma)(\epsilon+L_{f}^{\sigma}\sigma+\frac{\sqrt{s_{A}}\sigma^{2})^{2}}{2\gamma^{2}}}.\end{split}$ (9)
  • (ii)
    Suppose $\gamma$ is the penalty parameter such that
@@
Theorem 1 further indicates that selecting $\gamma=O(1/\sqrt{\epsilon})$ is requisite for achieving an $\epsilon$-KKT point of problem (4). As we will see later, this theoretical insight corroborates the empirical evidence observed in our study.
@@ 2.2 The ADMM Approach
By leveraging this bilinear factorization, (6) is rendered compatible with the ADMM approach shown in Algorithm 1, owing to the algorithm’s inherent ability to handle the variables $U$ and $V$ separately. Firstly, we define the augmented Lagrangian function of (6) as follows
$\mathcal{L}_{\rho}(U,V,{\lambda})=f(UV^{\top})+\frac{\gamma}{2}\|U-V\|_{F}^{2}+\big{\langle}{\lambda},\mathcal{A}(UV^{\top})-b\big{\rangle}+\frac{\rho}{2}\|\mathcal{A}(UV^{\top})-b\|^{2}.$ (10)
Since $f$ is convex, it is easy to see that both the subproblems of ADMM are $\gamma$-strongly convex. The optimality conditions of the subproblems can be written as
@@
$\begin{split}0=&\nabla f(V^{k+1}(U^{k+1})^{\top})U^{k+1}-\gamma(U^{k+1}-V^{k+1})+\sum_{i=1}^{m}\lambda_{i}^{k}A_{i}U^{k+1}\\ &+\rho\sum_{i=1}^{m}(\big{\langle}A_{i}U^{k+1},V^{k+1}\big{\rangle}-b_{i})A_{i}U^{k+1}.\end{split}$ (12)
For linear SDP instances, it is noteworthy to point out that because the $\gamma$ penalty term from the splitting, the subproblems (11) and (12) reduce to two well-conditioned linear systems of equations, which can be solved effectively using the conjugate gradient method. Detailed Numerical results supporting this idea can be found in Section 4.4.

### 25 Mar 2024 v2

@@ Abstract
We introduce a new first-order method for solving general semidefinite programming problems, based on the alternating direction method of multipliers (ADMM) and a matrix-splitting technique. Our algorithm has an advantage over the Burer-Monteiro approach as it only involves much easier quadratically regularized subproblems in each iteration. For a linear objective, the subproblems are well-conditioned quadratic programs that can be efficiently solved by the standard conjugate gradient method. We show that the ADMM algorithm achieves sublinear or linear convergence rates to the KKT solutions under different conditions. Building on this theoretical development, we present LoRADS, a new solver for linear SDP based on the Low-Rank ADMM Splitting approach. LoRADS incorporates several strategies that significantly increase its efficiency. Firstly, it initiates with a warm-start phase that uses the Burer-Monteiro approach. Moreover, motivated by the SDP low-rank theory (So et al., 2008), LoRADS chooses an initial rank of logarithmic order and then employs a dynamic approach to increase the rank. Numerical experiments indicate that LoRADS exhibits promising performance on various SDP problems. A noteworthy achievement of LoRADS is its successful solving of a matrix completion problem with $15,694,167$ constraints and a matrix variable of size $000\times 40,$ 000 × 40 , 000 in $351$ seconds.
@@ 1 Introduction
SDPs have a wide range of applications in practice, such as the MaxCut problem (Goemans and Williamson, 1995), optimal power flow (Lavaei and Low, 2011), combinatorial optimization (Boyd and Vandenberghe, 1997) and sensor network localization (So and Ye, 2007), to name a few. Traditionally, interior point methods are the go-to solution for solving SDP problems. It has been well known that for SDP, interior point methods can achieve high accuracy in polynomial time. See  Alizadeh (1995); Nesterov and Nemirovskii (1994); Alizadeh (1991). Despite their effectiveness, traditional interior point methods struggle with scalability for large-scale problems due to the necessity of solving Newton systems repeatedly, which is computationally intensive. To bypass the Newton system and enhance scalability, more recent research has shifted towards the first-order methods (Wen et al., 2010; Lan et al., 2011; O’donoghue et al., 2016). A significant area of interest has been the use of penalty methods, such as the augmented Lagrangian method and alternating direction method of multipliers (ADMM), for solving SDPs (Wen et al., 2010; Zhao et al., 2010; Yang et al., 2015; Wang et al., 2023). Specifically, Zhao et al. (2010) introduced SDPNAL, which employs the augmented Lagrangian method on dual SDP and uses an inexact semi-smooth Newton-CG method to solve the inner subproblem. The subsequent work (Yang et al., 2015) presented an enhanced algorithm SDPNAL+, which achieves improved efficiency and robustness and adequately addresses the numerical difficulty of SDPNAL when solving degenerate SDPs. O’donoghue et al. (2016) proposes a general splitting conic solver (SCS) for linear and conic programming, which applies ADMM to solve the homogeneous self-dual embedding associated with SDP. For large-scale SDP, a critical issue is the quadratically growing ($\mathcal{O}(n^{2})$) memory cost in maintaining the matrix variable. For a class of rank-constrained matrix problems, Yurtsever et al. (2017) apply the conditional gradient method to their convex relaxation problems. To reduce the storage overheads, they employ random sketching to keep track of a low-rank approximation of the matrix variable. In a follow-up work by Yurtsever et al. (2021), random sketching is further incorporated in the augmented Lagrangian-based method to deal with more general trace-constrained SDPs.
@@
Burer and Monteiro (2003, 2005) propose an effective limited memory BFGS augmented Lagrangian algorithm to handle the low-rank formulation. Another popular low-rank approach for SDP is the Riemannian optimization methods (Journée et al., 2010; Wen and Yin, 2013; Boumal, 2015). These methods are built upon the assumption that the feasible set of (3) is a Riemannian manifold. Within the low-rank formulation, Tang and Toh (2023) proposes a rank-support adaptive Riemannian optimization method for solving the general convex SDP problems and gives the convergence rate analysis. There are also several different ways to solve the low-rank formulation of SDP. Bellavia et al. (2021) propose a relaxed variant of the interior point method, which exploits the low-rank structure of the optimal solution. Erdogdu et al. (2022) apply the block-coordinate ascent method to solve the low-rank reformulation of SDP with diagonal constraints and establish global convergence towards the first-order stationary point. In the BM formulation, a critical step is determining the rank of the matrix variable. The rank $r$ is chosen to ensure that either an optimal solution for (3) is attainable or a satisfactory approximate low-rank solution is obtained. In particular, So et al. (2008) demonstrate the existence of a solution of logarithmic rank order that still maintains good approximation property. This discovery forms the basis of an empirical tactic to initiate the algorithm with a solution of rank $\mathcal{O}(\log(m))$, which will be employed in our approach.
@@
Our paper’s contributions are summarized as follows: First, we introduce a novel ADMM algorithm tailored for general SDPs that accommodate nonlinear and convex objectives. By introducing a new matrix splitting approach, we recast the SDP into a bilinearly constrained and quadratically penalized nonconvex problem, much different from the Burer-Monteiro approach. This formulation offers two primary benefits for ADMM implementation: 1) For a linear objective, enabled by the splitting technique, both the alternating steps in ADMM result in a convex and well-conditioned quadratic problem, and hence, can be efficiently solved by either direct methods or indirect ones such as the conjugate gradient descent method. This simplification provides a substantial advantage over the nonconvex quartic subproblems inherent in the BM formulation; 2) This splitting provides an enlarged search area, which could potentially improve the algorithm’s performance on some hard instances. Under mild conditions, we prove that the ADMM algorithm achieves a sublinear rate of convergence to the KKT solutions. Additionally, when the Łojasiewicz inequality is satisfied with an exponent of ${1}/{2}$, we enhance the convergence result to a linear rate, offering a significant improvement in the algorithm’s efficiency.
@@
Second, we present a new SDP solver, which employs low-rank factorization and ADMM through matrix splitting (LoRADS). This solver integrates the proposed ADMM along with several new implementation strategies that significantly improve the practical efficiency. Our empirical evaluations highlight two critical heuristics: 1) a warm-start procedure employing the Burer-Monteiro algorithm, and 2) aggressive rank reduction to logarithmic order with a dynamic rank selection strategy ensuring the stability. By leveraging these enhancements, we compare LoRADS with the leading SDP solvers on a wide range of problems, including MaxCut, matrix completion, and general SDP benchmarks in SDPLIB and Mittlemann’s dataset. Extensive numerical studies demonstrate that LoRADS is highly competitive and often outperforms the popular SDP solvers. Furthermore, LoRADS demonstrates notable efficacy in addressing super large-scale problems, an area where current SDP solvers show limitations. A case in point is LoRADS’s ability to solve a matrix completion problem with $15,694,167$ constraints and a $40000\times 40000$ matrix variable within approximately 350 seconds.
@@
The rest of our paper is organized as follows. Section 2 introduces the proposed ADMM algorithm with matrix splitting for semidefinite programming. Section 3 establishes the convergence and convergence rate of the proposed algorithm. The missing proof can be found in Appendix A. Section 4 develops an enhanced solving approach based on ADMM and several useful implementation heuristics. Section 5 examines the empirical performance of our solver on various SDP problems. Finally, Section 6 concludes our paper and points out our future direction.
@@
Let $\|\cdot\|_{2}$ denote the Euclidean norm for vectors and operator norm for matrices, and $\|\cdot\|_{F}$ denote the Frobenius norm for matrices. For a given matrix $X$ and radius $\delta$, we use $\mathcal{B}_{\delta}\big{(}X\big{)}$ to denote the ball of radius $\delta$ around $X$, i.e. $\mathcal{B}_{\delta}\big{(}X\big{)}=\{Y|\|X-Y\|_{F}\leq\delta\}$. For a nonempty and compact set $\mathcal{S}$, we reuse the notation $\mathcal{B}_{\delta}\big{(}\mathcal{S}\big{)}$ to denote the neighborhood of $\mathcal{S}$, which is defined as $\mathcal{B}_{\delta}\big{(}\mathcal{S}\big{)}=\{Y|\inf_{Z\in\mathcal{S}}\|Y-Z\|_{F}\leq\delta\}$. For any matrix $X$, let $\sigma_{\min}(X)$ denote the minimal singular value of $X$. Let $\|\mathcal{A}\|_{2}$ denote the operator norm of the linear map $\mathcal{A}$ and $s_{A}=\sum_{i=1}^{m}\|A_{i}\|_{F}^{2}$. Given a sequence of matrices $\{X^{k}\}_{k=0}^{K-1}$ and a sequence of vectors $\{x^{k}\}_{k=0}^{K-1}$, we use $\Delta X_{k}$ to denote $X^{k}-X^{k-1}$ and $\Delta x^{k}$ to denote $x^{k}-x^{k-1}$, respectively.
@@ 2.1 Bilinear Reformulations and The Penalized Problem
To address the challenges posed by large-scale SDP instances and to thoroughly explore the low-rank domain, we propose a different and bilinear factorization technique for problem (1), yielding the following formulation:
@@
To elucidate the relationship between model (4) and (6) with a suitably large parameter $\gamma$, we introduce the following assumption.
@@
where $\nabla^{2}f(X)[H]$ denotes the tensor product of $\nabla^{2}f(X)$ and $H$. From the symmetry of $f$, it is easy to verify that the partial gradients satisfy
$\frac{\partial{f(UV^{\top})}}{\partial U}=\nabla f(UV^{\top})V\quad{\rm and}\quad\frac{\partial{f(UV^{\top})}}{\partial V}=\nabla f(VU^{\top})U.$
@@
$\begin{split}\Big{\|}2\nabla f(U^{*}{(U^{*})}^{\top})U^{*}+\sum_{i=1}^{m}\lambda_{i}^{*}A_{i}U^{*}\Big{\|}_{F}\leq\epsilon,\\ \big{\|}\mathcal{A}(U^{*}{(U^{*})}^{\top})-{b}\big{\|}_{2}\leq\epsilon,\end{split}$ (7)
@@
$\begin{split}\Big{\|}\nabla f(\overline{U}\overline{V}^{\top})\overline{V}+\gamma(\overline{U}-\overline{V})+\sum_{i=1}^{m}\overline{\lambda}_{i}A_{i}\overline{V}\Big{\|}_{F}\leq\epsilon,\\ \Big{\|}\nabla f(\overline{V}\overline{U}^{\top})\overline{U}+\gamma(\overline{V}-\overline{U})+\sum_{i=1}^{m}\overline{\lambda}_{i}A_{i}\overline{U}\Big{\|}_{F}\leq\epsilon,\\ \big{\|}\mathcal{A}(\overline{U}\overline{V}^{\top})-{b}\big{\|}_{2}\leq\epsilon.\end{split}$ (8)
@@
Let $(\overline{U},\overline{V},\overline{{\lambda}})$ be any $\epsilon$-KKT point of problem (6) in $\mathcal{B}_{\sigma}\big{(}0^{n\times r},0^{n\times r},0^{m}\big{)}$ and $\widehat{U}=\frac{\overline{U}+\overline{V}}{2}$. Suppose Assumption 1 holds, then the following statement s valid.
$\begin{split}&\left\|{\cal A}(\widehat{U}{\widehat{U}}^{\top})-b\right\|_{2}\leq\epsilon+\frac{\|\mathcal{A}\|_{2}(\epsilon+L_{f}^{\sigma}\sigma+\sqrt{s_{A}}\sigma^{2})^{2}}{4\gamma^{2}},\\ &\left\|2\nabla{f(\widehat{U}\widehat{U}^{\top})}\widehat{U}+2\sum_{i=1}^{m}\overline{\lambda}_{i}A_{i}\widehat{U}\right\|_{F}\leq 2\epsilon+\frac{(2L_{0}^{\sigma}\sigma^{3}+3L^{\sigma}\sigma)(\epsilon+L_{f}^{\sigma}\sigma+\sqrt{s_{A}}\sigma^{2})^{2}}{2\gamma^{2}}.\end{split}$ (9)
Moreover, suppose that we choose the penalty parameter $\gamma$ such that
@@
Theorem 1 further indicates that if we have an $\epsilon$-KKT point of the penalty problem with $\gamma=\Omega(1/\sqrt{\epsilon})$, then we are able to obtain an $\epsilon$-KKT point of problem (4). As we will see later, this theoretical insight corroborates the empirical evidence observed in our study.
@@ 2.2 The ADMM Approach
We describe ADMM for solving the penalty problem (6) in Algorithm 1. In this algorithm, the augmented Lagrangian function of (6) is defined by
$\mathcal{L}_{\rho}(U,V,{\lambda})=f(UV^{\top})+\frac{\gamma}{2}\|U-V\|_{F}^{2}+\big{\langle}{\lambda},\mathcal{A}(UV^{\top})-b\big{\rangle}+\frac{\rho}{2}\|\mathcal{A}(UV^{\top})-b\|_{2}^{2}.$ (10)
By leveraging the bilinear factorization, ADMM alternates between optimizing two proximal subproblems associated with $U$ and $V$. Since $f$ is convex, both the subproblems are $\gamma$-strongly convex with their own variables. The optimality conditions of the subproblems can be written as
@@
$\begin{split}0=&\nabla f(V^{k+1}(U^{k+1})^{\top})U^{k+1}-\gamma(U^{k+1}-V^{k+1})+\sum_{i=1}^{m}\lambda_{i}^{k}A_{i}U^{k+1}\\ &+\rho\sum_{i=1}^{m}\big{(}\big{\langle}A_{i}U^{k+1},V^{k+1}\big{\rangle}-b_{i}\big{)}A_{i}U^{k+1}.\end{split}$ (12)
For linear SDPs, it is noteworthy to point out that due to the $\gamma$ penalty term from the splitting, both the subproblems (11) and (12) reduce to well-conditioned linear systems, which can be solved effectively by the conjugate gradient (CG) method. Numerical results supporting this view can be found in Section 4.4.